OpenStudy (anonymous):

Consider the two function machines that are placed as shown. Find the final output for 0,5,12. f(x)=8x g(x)=x+3 (g°f)(x) The output for 0 is... The output for 5 is... The output for 12 is...

OpenStudy (apoorvk):

(g*f)(x) = g(f(x)) so here, (g*f)(x) = (8x)+3 now plug in values of x, solve, and enjoy!

OpenStudy (anonymous):

\[g(f(x)) = g(8x) = 8x+3\] \[g(f(0)) = 8(0)+3=3\] \[(g(f(5))=8(5)+3=43\] \[g(f(12))=8(12)+3 = 99\]

OpenStudy (anonymous):

Thanks for showing me how it works. I can do my other problems like this one now.